HEELLP!!! here is the equation: 1/x^2-25 - 4/x-5 = 2/x+5 ANYBODY! PLEASE!!
Answer by stanbon(75887)

here is the equation: 1/x^2-25 - 4/x-5 = 2/x+5
---
Multiply thru by x^2-25 to get:
1 - 4(x+5) = 2(x-5)
1-4x-20 = 2x-10
-19-4x = 2x-10
6x = -9
x = -3/2
==================
Cheers,
Stan H.